If the measure of angle P of triangle PQR is 3x, the measure of angle Q is 5x and the measure of angle R is 4x. What is the value of x?
Answer by ikleyn(52781)

.

Since the sum of interior angles of any triangle is 180 degrees,

you have this equation to find x


    3x + 5x + 4x = 180  degrees


    12x          = 180 degrees


      x          = 180/12 = 15 degrees.      ANSWER